OpenStudy (anonymous):

Help A suitcase measures 28 inches long and 16 inches high. What is the diagonal length of the suitcase? Round your answer to the nearest tenth of an inch. A.33.5 in B.32.2 in C.12.8 in D.29.3 in

OpenStudy (anonymous):

@cwrw238 can you help me please?

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

what you can do in this problem is make up a triangle. since the suitcase is rectangular, draw a diagonal line and you will have two right triangles. now, get one triangle and plug in the information given. the length is 28 and the height is 16. now use the Pythagorean theorem to find out the hypotenuse

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\] \[28^{2}+16^{2}=c ^{2}\] 784 + 256 = c^2 1040=c^2 get rid of c squared by applying the square root to both sides, sqrt(1040)=c 32.249 = c

OpenStudy (anonymous):

@ivan1019 thankss